OpenStudy (anonymous):

Help me find these? Vertex, Focus, Directrix, and Length of focal chord. for y^2+6y-2x+13=0 Any help is appreciated!

OpenStudy (campbell_st):

complete the square in y after rewriting the equation as \[y^2 + 6y = 2x - 13 \]

OpenStudy (campbell_st):

this is a parabola which is concave right... the general form is \[(y - k)^2 = 4a(x - h)\] the vertex is (h, k) a = focal length

OpenStudy (anonymous):

so vertex would be -13,6 and focal length is 2? @campbell_st

OpenStudy (campbell_st):

Ummm no its \[(y + 3)^2 = 2(x - 2)\] so it looks like the focal length is 1/2 and the vertex is (2, -3)

OpenStudy (campbell_st):

your parabola looks like